# Generative Models and Connected and Automated Vehicles: A Survey in Exploring the Intersection of Transportation and AI

### v1

@@ II-A History of Generative Models
The development of Generative Models began with foundational work in AI and machine learning, including the LISP programming language in the 1960s, the ELIZA chatbot [64], and early expert systems like Dendral [65] and MYCIN [66]. The rise of the internet and advancements in computing power in the 1990s and 2000s led to significant progress in machine learning, neural networks, and deep learning, setting the stage for modern generative AI [68]. Figure 1 showcases a significant evolution from unimodal approaches in natural language processing (NLP) and computer vision (CV) towards increasingly sophisticated multimodal technologies. Early models like N-Gram [69] and GANs laid the groundwork between 2000 and 2015. The period from 2015 to 2018 saw the introduction of transformative architectures such as Transformers [70] and the emergence of multimodal models like StyleNet [71]. This evolution accelerated from 2018 to 2020 with advancements like BERT [72], GPT-2 [73], and StyleGAN, expanding to complex multimodal approaches including VisualBERT [74]. The trend from 2020 to 2023 highlights the proliferation of large language models like GPT-3 [75] and innovative visual technologies such as DALL-E [76]. From the year 2023 up until now, we have witnessed a steady rise in the emergence of innovative models like the remarkable GPT-4 [77] and the revolutionary Sora [78]. This trend signifies the continuous evolution and advancement within the field of technology and artificial intelligence.
@@ II-B Challenges in the Generative Models
Generative Adversarial Networks (GAN) [7] [8]
Realistic Data Generation: GANs can produce highly realistic synthetic data, aiding in diverse scenario training for automotive vehicles (AV) systems without costly real-world data collection.
@@
Reinforcement Learning [9] [10]
Adaptive Decision Making: RL models are excellent at learning optimal actions through trial and error, enabling autonomous vehicles to adapt to changing road conditions dynamically.
@@
StyleGAN [11][12]
High-Quality Images: Produces high-resolution, photo realistic images with fine details.
@@
Neural Architecture Search(NAS) [13][14]
Automation: NAS automates the design of network architectures, potentially outperforming manually designed networks, especially in multi-objective optimization scenarios.
@@
Collaborative AI [15][16][17]
Enhanced Learning and Adaptation: Collaborative AI allows vehicles to learn from each other’s experiences, significantly improving their ability to adapt to new environments and situations without direct human intervention.

### v2

@@ II-A History of Generative Models
The development of Generative Models began with foundational work in AI and machine learning, including the LISP programming language in the 1960s, the ELIZA chatbot [64], and early expert systems like Dendral [65] and MYCIN [66]. The rise of the internet and advancements in computing power in the 1990s and 2000s led to significant progress in machine learning, neural networks, and deep learning, setting the stage for modern generative AI [68]. Figure 1 showcases a significant evolution from unimodal approaches in natural language processing (NLP) and computer vision (CV) towards increasingly sophisticated multimodal technologies. Early models like N-Gram [69] and GANs laid the groundwork between 2000 and 2015. The period from 2015 to 2018 saw the introduction of transformative architectures such as Transformers [70] and the emergence of multimodal models like StyleNet [71]. This evolution accelerated from 2018 to 2020 with advancements like BERT [72], GPT-2 [73], and StyleGAN, expanding to complex multimodal approaches including VisualBERT [74]. The trend from 2020 to 2023 highlights the proliferation of large language models like GPT-3 [75] and innovative visual technologies such as DALL-E [76]. From 2023 onward, we have witnessed a steady rise of innovative models like the remarkable GPT-4 [77]. However, the core architectures of LLMs have largely stabilized, and now research attention has increasingly shifted toward the development of Multimodal Large Language Models (MLLMs). This includes the rise of large vision-language models [98, 99, 100] capable of processing image-text pairs, followed by more general “any-to-text” models that accept inputs across modalities—such as images, video, audio, and sensor data—and generate coherent text outputs [101]. The latest frontier involves “any-to-any” generation models that can flexibly take in and produce outputs across multiple modalities [102, 103, 104].
As both LLMs and MLLMs continue to mature, researchers are increasingly integrating them into agentic frameworks [105, 106], enabling these models to interact with external tools and environments. In this paradigm, LLMs and MLLMs act as the “brains”, capable of reasoning, planning, and language understanding, while the agentic system provides the “arms,” granting them the ability to perceive the world, execute actions, and call APIs or retrieve external knowledge. This combination marks a transformative shift from static model usage toward dynamic, goal-directed agents. For example, in the context of autonomous driving, these agents can interpret visual scenes, reason about traffic conditions, plan maneuvers, and take appropriate actions in real time. Models such as Sora [78], DriveGPT [107], and Genie [108] exemplify this integration by coupling generative understanding with action-oriented execution. This synergy between multimodal understanding and interactive agency is paving the way for next-generation intelligent systems that are both perceptive and capable of autonomous decision-making in complex real-world environments.
@@ II-B Challenges in the Generative Models
Generative Adversarial Networks (GAN) [7] [8] Realistic Data Generation: GANs can produce highly realistic synthetic data, aiding in diverse scenario training for automotive vehicles (AV) systems without costly real-world data collection.
@@
Reinforcement Learning [9] [10] Adaptive Decision Making: RL models are excellent at learning optimal actions through trial and error, enabling autonomous vehicles to adapt to changing road conditions dynamically.
@@
StyleGAN [11][12] High-Quality Images: Produces high-resolution, photo realistic images with fine details.
@@
Neural Architecture Search(NAS) [13][14] Automation: NAS automates the design of network architectures, potentially outperforming manually designed networks, especially in multi-objective optimization scenarios.
@@
Collaborative AI [15][16][17] Enhanced Learning and Adaptation: Collaborative AI allows vehicles to learn from each other’s experiences, significantly improving their ability to adapt to new environments and situations without direct human intervention.